# Cloud gap-filling with deep learning for improved grassland monitoring

### v1

@@ 3.2 Sentinel Data
In this study, we utilized both Sentinel-1 and Sentinel-2 satellite imagery. We obtained Sentinel-1 GRD and SLC, as well as Sentinel-2 L1C data for the year 2020 from the CreoDIAS data repository. It is worth noting that the data acquisition period is before December 2021, prior to the instrument electronics power supply anomaly occurred on Sentinel-1B (Pinheiro2022, ). As a result, the revisit time for the radar data is 6 days.
The study areas are spread evenly throughout the Lithuania, with multiple orbits intersecting with each other to cover the regions of interest. From the total of 14 available Sentinel-2 tiles, 8 (i.e., 34UEG, 34UFG, 34UFF, 34VEH, 35ULA, 35ULB, 35UMB and 35VLC) intersect with the respective study sites. Moreover, two Sentinel-1 relative orbits were selected, namely 58 and 131, which in combination cover the entire country of Lithuania and guarantees consistent acquisition geometry of the respective time series per study site. Additionally, these specific orbits are acquired in the afternoon, which is crucial for grassland monitoring since morning dew can cause unexpected decrease of coherence on both mowed and unmowed parcels (tamm2016, ).
@@ 3.2.1 Data Processing
A time series of Sentinel-2 images was collected between April 1st and September 30th, 2020. The images were atmospherically corrected using the sen2Cor software (sen2cor_new, ), transforming them from Top of Atmosphere (TOA) Level 1C to Bottom of Atmosphere (BOA) Level 2A products. The resulting products also contain the Scene Classification Layer (SCL), which includes a pixel classification map for clouds, cloud shadows, vegetation, and other features. The images were then reprojected and clipped to the areas of interest. Our focus is primarily on the calculation of NDVI (see Table 2), which uses the B04 and B08 bands representing the red and near-infrared wavelengths, respectively. Both bands have a spatial resolution of 10 m.
We used Level-1 Ground-Range-Detected (GRD) products from Sentinel-1 imagery in Interferometric Wide (IW) swath mode to generate backscatter coefficient ($\mathrm{\sigma_{0}}$) and Single-Look-Complex (SLC) products, which were used to compute the interferometric coherence coefficients. The GRD data pre-processing was performed using the snappy library, which included clipping to the area of interest, radiometric calibration, speckle filtering with the Refined Lee Filter (RLF), terrain correction using Shuttle Radar Topography Mission (SRTM) 10-m, and conversion of backscatter coefficient into decibels (dB). For the SLC products, we applied TOPS Split to select only the essential bursts for analysis. The next step was to retrieve and apply orbit auxiliary data that contained information on the satellite’s position during SAR data acquisition. Then, the Sentinel-1 Back-Geocoding process was used to co-register the split products based on updated orbit information and a Digital Elevation Model (DEM). Finally, we estimated the coherence between the reference and secondary images, taking into account any differences in imaging conditions, which provided information on any changes or movements that may have occurred between the two acquisitions. All coefficients were computed in VV and VH polarization modes.
Furthermore, apart from the backscatter and coherence coefficients (in VV and VH polarization), we also calculated four products based on combinations of them, namely the mixed coherence, the Radar Vegetation Index (RVI), the backscatter ratio and the backscatter cross-ratio. In Table 2, we summarize the list of all features extracted from Sentinel-1 and Sentinel-2 for the development of this study.
@@
To handle the large amount of data efficiently, we utilized the Agriculture monitoring Data Cube (ADC) developed by datacube_2022 . The acquisition dates for the Sentinel satellites varied by region, so we brought all data components into a common temporal framework based on the acquisition dates of Sentinel-1 data, in each area of interest, to ensure consistency in our dataset. Since most regions were covered by relative orbit 131, we used its acquisition dates as a reference for the common temporal framework. We then shifted all other measurements from Sentinel-2 and Sentinel-1 (orbit 58) satellites to the closest Sentinel-1 date of the 131 orbit. As a result, the final dataset consists of time series with a 6-day time step from April 9th, 2020 until September 24th, 2020, containing 29 individual timestamps. We shifted Sentinel-2 cloud-free timestamps to the closest date of reference only if there was a 3-day distance between them. Otherwise, we denoted the corresponding timestamps as empty (i.e., with NaN values).
We used the same interpolation method to fill missing values for pixels that had been masked out due to clouds from the SCL product of the sen2Cor software. However, these cloud masks are not always accurate, especially in cases of cloud shadows, thin clouds, or pixels adjacent to clouds, which cause an underestimation of the NDVI value that can potentially lead to incorrect detection of mowing events (tarrio2020, ). To address this issue, we implemented Algorithm 1, which is a threshold-based module, and applied it to each pixel’s time series. The aim of this module is to remove cases of a sudden drop in NDVI (parameter $\mathrm{\alpha}$) followed by a relevant sudden increase (parameter $\mathrm{\beta}$) between two successive time-steps (Fig. 2). In our case, we configured both $\mathrm{\alpha}$ and $\mathrm{\beta}$ threshold parameters for the sudden drop and relevant increase to 0.15. This deliberate choice of a relatively high threshold, minimizes the overall number of cases that are removed, ensuring that only certain noisy timestamps are filtered out. Additionally, we examined the difference between the NDVI values in the timestamps before and after the examined one, to prevent removal of cases caused by gradual NDVI decline due to grassland senescence. This difference should be at least -0.05 (parameter $\mathrm{\gamma}$). An equivalent approach has been developed by komisarenko2022 using NDVI triplets and the rhombus equation for the identification of outlier cases. Similarly, in the recent study of WATZIG2023113577 a gradient boosting algorithm is trained to detect false positive mowing predictions caused by missed clouds.
@@
Figure 2: Example of NDVI time series outlier removal. The red dashed line shows the raw NDVI time series (i.e., including noise), and the green line illustrates the NDVI after filtering out potential cloudy measurements that remained following the cloud masking step.
@@ 3.2.2 Annotated data
To streamline the assessment process, we utilized validation data from the Lithuanian Paying Agency (NPA) as a starting point within the regions of interest. The parcels selected were compliant with national mowing regulations, meaning that at least one agricultural event had occurred by August 1st, although the exact timing of this event was unknown. Moreover, for the photo-interpretation process we selected parcels with minimal cloud coverage during the inference period. This approach minimized the risk of overlooking cases that might have gone undetected due to persistent cloud cover, as no extended gaps beyond a few days were observed in these timestamps.
The photo-interpretation process was carried out by three experts that used multi-temporal Sentinel-1 and Sentinel-2 images as reference. Since Sentinel satellites do not offer daily observations, the labeling process involved identifying the closest timestamps before and after a grassland event took place, from early April to late September. Each expert independently assigned labels on the acquisition dates of Sentinel-2, indicating the date before and after a grassland event has occurred. The samples were distributed evenly across the six sites to ensure reliable model optimization and results assessment. Since SAR coefficients are sensitive to changes in the canopy of grasslands, the photo-interpretation analysis additionally incorporates Sentinel-1 data. Steep changes between successive images may showcase the occurrence of a potential grassland event. For this reason, true color and false color (B08-B04-B03) composites of Sentinel-2 time series have been found to be particularly useful in identifying the presence of an event (Fig. 3).
@@
Then, the examination of NDVI and SAR multi-temporal signatures were used to guide the decision of the photo-interpreter. Their decision was dictated by the fact that NDVI gets much lower values after an event (at least a decrease of 0.1), while SAR coefficients acts the opposite (Rules 1-2). Secondly, an event must take place on compact dense areas and not scattered spots inside grassland parcels (Rule 3). If all the above are met, the final decision lies on the consequent values of NDVI (Rule 4). Cases where NDVI does not increase shortly (usually up to 15 days) after the potential event are excluded and are not considered as valid events. This behaviour many times resemble harvest events, and thus such cases may have been wrongly characterized as grasslands.
@@
The procedure was applied in all six study sites and the final dataset includes 803 parcels, which correspond to 287,281 individual pixels (see Table 3). In this dataset, 650 cases had one mowing event, 99 cases multiple mowing events, and 54 cases showed no evidence of mowing. Notably, a significant amount of the identified mowing events occurred during the June and early July (refer on Fig. S2 in the A for details).
@@
Region # parcels # pixels Region 1 274 64,141 Region 2 189 106,556 Region 3 98 33,065 Region 4 91 39,254 Region 5 79 23,639 Region 6 72 20,626 Total 803 287,281
@@ 4.1 Cloud coverage distribution
The study regions are characterized by unique cloud coverage profiles. Fig. 4 visualizes the percentage of grassland parcels affected by clouds on each date and each study region. We observe significantly lower cloud coverage during summer months (from June to August), when most of mowing events occur. However, there are still cases that grassland activity is not observed or not identified soon enough due to clouds.
@@
Figure 4: The percentage of grassland parcels affected by clouds on each reference date. The colors in this figure correspond to the colors used for each study area, as these are depicted in Fig. 1.
@@
We determined the cloud coverage for each grassland parcel by calculating the ratio of cloudy Sentinel-2 observations to the total of acquisition dates. The mean cloud coverage for the entire dataset is approximately 45%. To ensure higher quality ground truth information, we worked with pixel time series and selected only the densest ones. Specifically, we excluded cases that had i) more than one successive missing value during June and July, ii) more than 3 missing values in total during June and July, iii) more than two consecutive missing values during the rest of the assessed period, and iv) an overall cloud coverage higher than 35%. This approach helps ensure that we did not omit any mowing events. Finally, given that in dense time series simple temporal interpolation methodologies are able to effectively fill the short-term gaps, we used the Akima method (akima1970, ) to generate the full target NDVI time series.
4.2 Artificial masking
@@
To train and validate our methodology, we used artificially generated cloud masks, which were used to mask out specific timestamps in the dense NDVI time series. For each region, we collected a set of binary time series indicating the presence or absence of clouds, based on the actual cloud coverage distribution in the corresponding area. Subsequently, we used a bootstrapping procedure to assign a random cloud mask to each NDVI time series. Further, we ensured consistency by applying the same mask to all pixels within the same field boundaries. This approach ensured standardization across various parcels and maintained the alignment with the typical cloud coverage distribution in each region. An example of the NDVI values of a pixel after applying a cloud mask is shown in Fig. 5.
@@
Figure 5: NDVI time series with artificial cloud masking. The green dots on top of the plot show the complete NDVI time series for a pixel post-processing, with the black and white dots directly beneath representing a simulated cloud mask. In the scatter plot, the NDVI time series after the application of the artificial mask is displayed. Non-filled green dots indicate values that have been masked and are treated as missing data, while the filled green dots denote the ground truth.
@@ 4.3 Model training and architecture
The proposed deep learning Sentinel-1/2 Fusion (SF) methodology is composed of sequential CNN and LSTM blocks (Fig. 6). It is a supervised-learning neural network architecture that aims to extract full NDVI time series from Sentinel-1 time series and the available cloud-free Sentinel-2 observations. This architecture is based on the one proposed by zhao2020 , with the main difference being the inclusion of masked NDVI time series in the input feature space.
Figure 6: Overview of the model’s architecture for reconstructing continuous NDVI time series with a fixed 6-day time step using SAR data and the available NDVI values.
@@ 4.3.1 CNN block
‘The CNN block (CB) is used to reduce the impact of noise in the input and extract meaningful features for the subsequent layers. The block consists of N parallel and independent between them, one-dimensional convolutional branches (Conv1D), where N is the number of input channels. Each Conv1D branch is followed by max-pooling layers to reduce the dimensionality of the output. Following this, a non-linear Rectified Linear Unit (ReLU) activation function has been applied to transform negative values into zeros. The resulting feature maps are passed to a fully-connected layer to provide the final embeddings into the subsequent LSTM blocks. In the case of NDVI branch, a masking layer is added as the first layer in order to filter out masked NDVI inputs, which account for cloud presence. This strategy ensures that the network considers the cloudy NDVI timestamps as hidden and tries to reconstruct the complete NDVI time series based on the available cloud-free observations and the additional SAR data. The output of each Conv1D branch is a collection of feature maps that represent various characteristics of the input data. These feature maps are subsequently concatenated along the channel dimension and then forwarded to the encoder-decoder block.
@@ 4.3.2 Encoder-decoder block
RNNs are commonly used to analyze sequential data, as they can effectively identify patterns and dependencies between consecutive time steps. One of their key benefits is the capacity to retain information from previous time steps and use it to predict the current state, while sharing parameters across all time steps. To handle variable-length inputs and outputs, we have employed an encoder-decoder architecture with Bi-directional LSTM modules. These modules process input sequences in both forward and backward directions, allowing them to capture context from past and future time steps. The final NDVI time series is produced by a fully-connected layer that receives the output of the encoder-decoder block (EDB). The EDB consists of LSTM modules that perform the following operations:
@@
where $\mathrm{x_{t}}$ is the input at time $\mathrm{t}$, $\mathrm{h_{t-1}}$ is the previous hidden state, $\mathrm{c_{t-1}}$ is the previous cell state, $\mathrm{i_{t}}$ is the input gate, $f_{t}$ is the forget gate, $\mathrm{o_{t}}$ is the output gate, $\mathrm{c_{t}}$ is the cell state at time $\mathrm{t}$, and $\mathrm{h_{t}}$ is the hidden state at time $\mathrm{t}$. The weights and biases of the LSTM are represented by $\mathrm{W}$ and $\mathrm{b}$, respectively. The sigmoid function is denoted by $\mathrm{\sigma}$, and the hyperbolic tangent function by $\mathrm{\tanh}$.

### v2

@@ 3.2 Sentinel Data
For the present study, we utilized both Sentinel-1 and Sentinel-2 satellite imagery. We obtained Sentinel-1 Ground-Range-Detected (GRD) and Single-Look-Complex (SLC) products, as well as Sentinel-2 L1C data for the year 2020 from the CreoDIAS data repository. It is worth noting that the data acquisition period is before December 2021, prior to the instrument electronics power supply anomaly that occurred on Sentinel-1B [63]. As a result, the revisit time for the radar data is approximately 6 days.
The study regions are spread evenly throughout Lithuania, with multiple orbits intersecting with each other to cover the regions of interest. From the total of 14 available Sentinel-2 tiles, 8 (i.e., 34UEG, 34UFG, 34UFF, 34VEH, 35ULA, 35ULB, 35UMB and 35VLC) intersect with the respective study regions. Moreover, two Sentinel-1 relative orbits were selected, namely 58 and 131, which in combination cover the entire country of Lithuania and guarantee consistent acquisition geometry of the respective time series per study region. Additionally, these specific orbits were acquired in the afternoon, which is crucial for grassland monitoring since morning dew can cause an unexpected decrease of coherence on both mowed and unmowed parcels [53].
@@ 3.2.1 Data Processing
A time series of Sentinel-2 images from April 1st to September 30th, 2020, were atmospherically corrected with Sen2Cor [64] to transform the Top of Atmosphere (TOA) Level 1C data into Bottom of Atmosphere (BOA) Level 2A products, also generating a Scene Classification Layer (SCL) with cloud and land feature classifications. The images were reprojected and clipped to the study areas, focusing on NDVI calculation from the B04 (red) and B08 (near-infrared) bands, both at 10 m resolution.
Sentinel-1 Level-1 GRD data in Interferometric Wide (IW) swath mode and SLC products were used to derive backscatter ($\mathrm{\sigma_{0}}$) and interferometric coherence coefficients, respectively. GRD products pre-processing with the snappy library included radiometric calibration, speckle filtering with the Refined Lee Filter (RLF), terrain correction using Shuttle Radar Topography Mission (SRTM) 10-m, and conversion of backscatter coefficients into decibels (dB). Selected bursts using the TOPS SPLIT function from SLC products were co-registered using Back-Geocoding with updated orbit information and a Digital Elevation Model (DEM) to align split images, allowing for coherence estimation between reference and secondary images. Finally, backscatter and coherence coefficients were generated in VV and VH polarization modes, along with derived products including mixed coherence, Radar Vegetation Index (RVI), backscatter ratio, and cross-ratio.
In Table 2, we summarize the list of all features extracted from Sentinel-1 and Sentinel-2 for the development of this study.
@@
To handle the large amount of data efficiently, we utilized the Agriculture monitoring Data Cube (ADC) developed by [65]. To ensure temporal consistency across regions, we established a common framework based on acquisition dates from Sentinel-1’s relative orbit 131 as the primary reference. This approach aligned Sentinel-1 and Sentinel-2 data to the closest dates on a unified 6-day interval from April 9 to September 24, 2020, resulting in 29 consistent timestamps across the study period. Cloud-free Sentinel-2 timestamps were shifted to match the nearest reference date if within a 3-day window; otherwise, gaps were marked as empty (i.e., with NaN values).
Missing values resulted from noisy pixels that had been masked out due to clouds from the SCL product of the sen2Cor software. However, these cloud masks are not always accurate, especially in cases of cloud shadows, thin clouds, or pixels adjacent to clouds, which cause an underestimation of the NDVI value that can potentially lead to incorrect detection of mowing events [66]. To address this issue, we implemented a threshold-based outlier removal algorithm (Algorithm 1) to filter NDVI values with sudden drops (threshold $\mathrm{\alpha}$) followed by a quick recovery (threshold $\mathrm{\beta}$) between two successive steps (Fig. 2). In our case, we configured both $\mathrm{\alpha}$ and $\mathrm{\beta}$ threshold parameters to 0.15. This deliberate choice of a relatively high threshold, minimizes the overall number of cases that are removed, ensuring that only certain noisy timestamps are filtered out. Additionally, to prevent filtering cases of gradual NDVI decline due to grassland senescence, an additional threshold $\mathrm{\gamma=-0.05}$ was used to compare neighboring values. An equivalent approach has been developed by [58] using NDVI triplets and the rhombus equation for the identification of outlier cases. Similarly, in the recent study of [67] a gradient boosting algorithm was trained to detect false positive mowing predictions caused by missed clouds.
@@
Figure 2: Example of NDVI time series outlier removal. The red dashed line shows the raw NDVI time series (i.e., including noise), and the green line illustrates the NDVI after filtering potential cloudy measurements remaining from the initial cloud masking step.
@@ 3.2.2 Annotated data
To streamline the assessment process, we utilized validation data from the Lithuanian Paying Agency (NPA) as a starting point within the regions of interest. The parcels selected were compliant with national mowing regulations, meaning that at least one agricultural event had occurred by August 1st, although the exact timing of this event was unknown. Moreover, an equal number of random samples were included. For the photo-interpretation process, we selected parcels with minimal cloud coverage during the inference period. This approach minimized the risk of overlooking cases that might have gone undetected due to persistent cloud cover.
The photo-interpretation process was carried out by three experts who used multi-temporal Sentinel-1 and Sentinel-2 images as reference. Since Sentinel satellites do not offer daily observations, the labeling process involved identifying the closest timestamps before and after a grassland event took place, from early April to late September. Each expert independently assigned labels on the acquisition dates of Sentinel-2, indicating the date before and after a grassland event occurred. The samples were distributed evenly across the six regions to ensure reliable model optimization and results assessment. Since SAR coefficients are sensitive to changes in the canopy of grasslands, the photo-interpretation analysis additionally incorporated Sentinel-1 data. Steep changes between successive images may showcase the occurrence of a potential grassland event. For this reason, true color (B04-B03-B02) and false color (B08-B04-B03) composites of Sentinel-2 satellite image time series are particularly useful in identifying the presence of an event (Fig. 3).
@@
Then, the examination of NDVI and SAR multi-temporal signatures was used to guide the decision of the photo-interpreter. Their decision was dictated by the fact that NDVI gets much lower values after an event (at least a decrease of 10%), while SAR coefficients act the opposite (Rules 1-2). Secondly, an event must take place in compact dense areas and not scattered spots inside grassland parcels (Rule 3). If all the above are met, the final decision lies on the consequent values of NDVI (Rule 4). Cases where NDVI does not increase shortly (usually up to 15 days) after the potential event are excluded and are not considered valid events. This behavior many times resembles harvest events, and thus such cases may have been wrongly characterized as grasslands.
@@
The procedure was applied in all six study regions and the final dataset includes 803 parcels, which correspond to 287,281 individual pixels (see Table 3). In this dataset, 650 cases had one mowing event, 99 cases had multiple mowing events, and 54 cases showed no evidence of mowing. Notably, a significant amount of the identified mowing events occurred during June and early July. Further analysis of the NDVI profiles revealed that larger drops between consecutive NDVI values were more prevalent in June and July, whereas August exhibited smaller differences, complicating the detection of mowing events. Moreover, the distribution of NDVI differences concerning the number of mowing events indicated significantly higher values for cases with a single mowing event. Conversely, cases with two mowing events showed smaller NDVI differences, especially for the second mowing event, suggesting a reduced impact on the NDVI profile (refer to Fig. S8 and S9 in the A for more details). This may be attributed to the prevalence of grazing practices in Lithuania [55].
@@
Region # parcels # pixels Region 1 (Alytus) 274 64,141 Region 2 (Klaipėda) 189 106,556 Region 3 (Utena) 98 33,065 Region 4 (Telšiai) 91 39,254 Region 5 (Panevėžys) 79 23,639 Region 6 (Šiauliai) 72 20,626 Total 803 287,281
@@ 4.1 Cloud coverage distribution
The study regions are characterized by unique cloud coverage profiles. Fig. 4 visualizes the percentage of grassland parcels affected by clouds on each date and each study region (see also the respective Fig. S2 in A). We observe significantly lower cloud coverage during the summer months (from June to August) when most mowing events occur. However, there are still cases of grassland activity that are not observed or not identified soon enough due to clouds.
@@
Figure 4: The percentage of grassland parcels affected by clouds on each reference date, for each region and overall.
@@
We determined the cloud coverage for each grassland parcel by calculating the ratio of cloudy Sentinel-2 observations to the total acquisition dates. The mean cloud coverage for the entire dataset is approximately 45%. To establish higher-quality ground truth information for the experiments, we worked with pixel time series and selected only the densest ones. Specifically, we excluded cases that had i) more than one successive missing value during June and July, ii) more than 3 missing values in total during June and July, iii) more than two consecutive missing values during the rest of the assessed period, and iv) an overall cloud coverage higher than 35%. This approach helped us ensure that we did not omit any mowing events. Finally, given that in dense time series simple temporal interpolation methodologies can effectively fill the short-term gaps, we used the Akima method [68] to generate the complete target NDVI time series (5339 parcels in total) served as ground truth.
4.2 Artificial masking and NDVI input component
@@
To train and validate our methodology, we employed artificially generated cloud masks to simulate real-world cloud coverage conditions in the NDVI time series data used as input. These cloud masks were applied to strategically obscure specific timestamps in the complete NDVI time series. For each region, we collected a set of binary time series indicating the presence or absence of clouds, based on the actual cloud coverage distribution in the corresponding area (see Section 4.1). Subsequently, we used a bootstrapping procedure to assign a random cloud mask to each NDVI time series of the same region. Fig. S3 (refer to section A) illustrates the comparison between the actual and artificial cloud coverage distributions for each region. Further, we maintained parcel consistency by applying the same mask to all pixels within the same field boundaries. Therefore, this approach ensured standardization across various parcels and maintained the alignment with the typical cloud coverage distribution in each region. An example of the NDVI values of a pixel after applying a cloud mask is shown in Fig. 5. Finally, input NDVI values affected by the generated cloud masks (i.e., artificial missing data) were uniformly replaced with a standardized placeholder value (e.g., -100) before being fed into the model.
@@
Figure 5: NDVI time series with artificial cloud masking. The green dots on top of the plot show the complete NDVI time series for pixel post-processing, where the black and white dots beneath represent a simulated cloud mask. In the plot, the NDVI time series after the application of the artificial mask is displayed. Non-filled green dots indicate values that have been masked and are treated as missing data (these were uniformly substituted with a standardized placeholder value before being inserted as input into the model), while the filled green dots denote the partially available data.
@@ 4.3 Model training and architecture
The proposed deep learning Sentinel-1/2 Fusion (SF) methodology is composed of sequential CNN and LSTM blocks (Fig. 6). It is a supervised-learning neural network architecture that aims to extract full NDVI time series from Sentinel-1 time series and the available cloud-free Sentinel-2 observations. This architecture is based on the one proposed by [29], with the key modification being the inclusion of masked NDVI time series in the input feature space.
Figure 6: An overview of the model architecture used for reconstructing continuous NDVI time series with a fixed 6-day interval, utilizing both SAR data and available NDVI values.
@@ 4.3.1 CNN block
The CNN block (CB) reduces the impact of noise from the input data and extracts meaningful features for the subsequent layers [29]. The block consists of N parallels and independent between them, one-dimensional convolutional branches (Conv1D), where N is the number of input channels. Each Conv1D branch is followed by max-pooling layers to reduce the dimensionality of the output. Following this, a non-linear Rectified Linear Unit (ReLU) activation function is applied to convert negative values to zeros. The resulting feature maps are passed to a fully connected layer to provide the final embeddings into the subsequent LSTM blocks. In the case of the NDVI branch, a masking layer is added as the first layer to filter out masked NDVI inputs (e.g., -100), ensuring that the model treats cloudy data points as hidden. This enables the network to focus on reconstructing missing NDVI values based on the available cloud-free observations and SAR data, effectively handling the sparsity introduced by cloud cover. The CNN block extracts temporal features from both the NDVI and SAR data, producing feature maps that capture various aspects of the input sequence. These maps are concatenated along the channel dimension before being passed into the LSTM encoder-decoder block, which handles the temporal dependencies in the data.
@@ 4.3.2 Encoder-decoder block
The encoder-decoder block (EDB), equipped with Bi-directional LSTM (BiLSTM) modules, is key to modeling the temporal dependencies within the input and the target NDVI time series. An LSTM is a type of RNN designed to capture both short-term and long-term dependencies by maintaining an internal memory state across time steps. This makes LSTMs particularly well-suited for sequential data, where past observations influence future predictions. In the EDB, the BiLSTM modules process input sequences in both forward and backward directions, enabling the model to capture context from both past and future time steps. The encoder part processes the input embedding sequence from the CB to generate hidden states, while the decoder uses these hidden states to generate the final predicted NDVI time series without missing values. The temporal dependencies between consecutive time steps are modeled as follows:
@@
where $\mathrm{x_{t}}$ is the input at time $\mathrm{t}$, $\mathrm{h_{t-1}}$ is the previous hidden state, $\mathrm{c_{t-1}}$ is the previous cell state, $\mathrm{i_{t}}$ is the input gate, $f_{t}$ is the forget gate, $\mathrm{o_{t}}$ is the output gate, $\mathrm{c_{t}}$ is the cell state at time $\mathrm{t}$, and $\mathrm{h_{t}}$ is the hidden state at time $\mathrm{t}$. The weights and biases of the LSTM are represented by $\mathrm{W}$ and $\mathrm{b}$, respectively. The sigmoid function is denoted by $\mathrm{\sigma}$, and the hyperbolic tangent function by $\mathrm{\tanh}$. The LSTM’s gating mechanisms (input, forget, and output gates) help manage information flow across time steps, ensuring that long-term dependencies are retained while mitigating the risk of vanishing gradients. The final NDVI time series is generated by passing the output of the EDB through a fully connected layer, which produces the reconstructed NDVI sequence based on the learned temporal patterns.